Solve.
9x^4-9x^2+2=0

Answer by atique.ah(27)

Solution
Let x^2=p therefore x^4=p^2
by substituting in the given problem we have
9p^2-9p+2=0
by breaking the middle term the equation can be written as
9p^2 -6p -3p +2=0
taking common
3p(3p-2)-1(3p-2)=0
(3p-1)(3p-2)=0
either 3p-1=0 or 3p-2=0
if3p-1=0 then p=1/3, if 3p-2=0 then p=2/3
by substituting x^=p
we have x=sqrt1/3 or sqrt2/3